OpenStudy (anonymous):

Determine the second and third terms of an arithmetic sequence if A)the first term is 6 and the 4th term is 33 Thanks)

OpenStudy (strawberry17):

It is going up by 9 :)

Directrix (directrix):

The two terms given are the fourth terms and the first terms. The fourth term = the first term + 3*d which is the common difference 33 = 6 + 3d d = 9 as @strawberry17 determined by inspection which is releatively easy when the common difference is an integer.

Directrix (directrix):

If the first term is 6, then then the second term will be 6 + 9 and the third term will be 6 + 2*9. The fourth term should be 6 + 9*3. If that cranks out to be 33, then the above work is correct. So, @noname81 what are the the second and third terms of the sequence? Post, and we will check them, okay?

OpenStudy (anonymous):

Im am trying to get use to the Tn=t1+(n-3)d formula I could have used logic to solve the equation but I would like to use this formula.

OpenStudy (anonymous):

I think that the 33=6+d will help. I can punch in the numbers into the equation. D=9 T1=6 n=3. Does that look about right ?

Directrix (directrix):

Okay, let's use that formula. Hold on.

Directrix (directrix):

The fourth term = 33 The first term = 6 Since t4 and t1 are (four - one = 3) places apart, then I know from the definition of an arithmetic sequence that t4 = t1 + 3*d. Using this, I can then solve for the common difference d. 33 = 6 + 3*d Solve for d. The common difference is 9. Do you agree @noname81 Questions?

OpenStudy (anonymous):

got it

Directrix (directrix):

The n-th term of an arithmetic sequence is of the form tn = t1 + (n – 1)d. Find t2 in the posted problem. t2 = t1 + (2-1)*9 = 6 + 1*9 = 15 for the second term.

OpenStudy (anonymous):

awsome crystal clear now. Thank you :)

Directrix (directrix):

Third Term: t3 = t1 + (3-1)*9 = 6 + 2*9 = 6 + 18 = 24 --> See if you agree @noname81 I apologize for my "false starts." But, I wanted to get the terminology right so that you can work all of these types of problems.

Directrix (directrix):

Here's a link to a great study resource: http://www.purplemath.com/modules/series.htm On the last two pages of that section are worked out solutions. Helpful.

OpenStudy (anonymous):

thanks

Directrix (directrix):

Glad to help.